Find the midpoint and distance of the line segment that connects the following points a) (0,3) and (4,7) b) (-9,6) and (-1, -2) c) (-4.5, -12.5) and (4.5, 13)


1
Expert's answer
2021-08-01T09:37:18-0400

The mid point of line segment joining two points ( x1, y1 ) and ( x2, y2 ) is given by


"(x, y) = \dfrac{x_1+x_2}{2}, \dfrac{y_1+y_2}{2}"



and the distance between two points is given by


"d = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}"



a) (0,3) and (4,7)


mid point "(x, y) = \dfrac{0+4}{2}, \dfrac{3+7}{2}"



mid point "(x, y) = (2,\ 5)"




"d = \sqrt{(4-0)^2+(7-3)^2}"


"d = 4\sqrt{2} \ units"





b) (-9,6) and (-1, -2)


mid point "(x, y) = \dfrac{-9-1}{2}, \dfrac{6-2}{2}"



mid point "(x, y) = (-5,\ 2)"




"d = \sqrt{(-1+9)^2+(-2-6)^2}"


"d = 8\sqrt{2} \ units"









c) (-4.5, -12.5) and (4.5, 13)


mid point "(x, y) = \dfrac{-4.5+4.5}{2}, \dfrac{-12.5+13}{2}"



mid point "(x, y) = (0,\ 0.25)"




"d = \sqrt{(4.5+4.5)^2+(13+12.5)^2}"


"d = 27.04 \ units"